OpenStudy (anonymous):

find the fourier series representation of f(x)=x + 1 for -1less than or equal to x less than or equal to 1

OpenStudy (tkhunny):

Have you considered doing the work? \(a_{0} = \int\limits_{-1}^{1}(x+1)\;dx = 2\) With that transformation, it magically turns into an Odd Function about y = 1, so we can forget about all the rest of the \(a_{i}\). \(b_{1} = \int\limits_{-1}^{1}(x+1)\cdot\sin(\pi\;x)\;dx\;=\;\dfrac{2}{\pi}\). Likewise, \(b_{2} = \int\limits_{-1}^{1}(x+1)\cdot\sin(2\pi\;x)\;dx\;=\;-\dfrac{1}{\pi}\). You're almost done!

OpenStudy (anonymous):

so does it mean i should the subtitute the result into f(x)=1ao/2 + sum( an cos nx t bn sinnx ) please can you complete the solving for me please

OpenStudy (tkhunny):

Yes, that is the right formula, however: 1) Pay attention. All the \(a_{n} = 0\), excepting the first. 2) Feel free to do your own integrals. You'll see a pattern, soon enough.

OpenStudy (anonymous):

i got f(x) = 1+1cosxsinx/pi -2cosxsinx/pi........ am i correct . if not please help with the solution

OpenStudy (tkhunny):

Not quite. cos(x) is incorrect in both. There is other stuff in the argument.